# Low-dimensional tori in Calogero–Moser–Sutherland systems

### v1

@@ Abstract.
We give an explicit description of low-dimensional Liouville tori in Calogero–Moser–Sutherland systems.
@@ 1. Introduction
Recall that Calogero–Moser–Sutherland system [1, 6, 8] is a model of $n$ one-dimensional particles on a circle with trigonometric potential. The phase space of this model is $T^{*}H_{\rm reg}$ with coordinates $(p_{i},\gamma_{i})$, $p_{i}\in\mathbb{R}$, $\lvert\gamma_{i}\rvert=1$, $\sum_{i=1}^{n}p_{i}=0$, $\prod_{i=1}^{n}\gamma_{i}=1$.
The symplectic structure is standard with Poisson brackets:
@@
(2) $H_{2}=\frac{1}{2}\sum_{i=1}^{n}p_{i}^{2}+\frac{1}{4}\sum_{i\neq j}^{n}\frac{c^{2}}{\sin^{2}(q_{i}-q_{j})}.$
@@
where $x$ is given by formula (12). Function $H_{k}$ are invariant with respect to simultenious permutations of $p$ and $q$. Thus, effective phase space is an orbifold $T^{*}H_{reg}/S_{n}$, where $S_{n}$ acts on $\mathfrak{h}^{*}\times H_{reg}$ diagonally.
The main result of this paper is the following explicit description of all Liouville tori. Let $(x_{1},\dots,x_{n})$ be the eigenvalues of $x$ (given (12)). Then:
  • when $x_{i}-x_{i+1}>c$ for all $i$, the level surface of $\left\{H_{k}\right\}_{k=2}^{n}$ is $(n-1)$-dimensional and with well-known angle coordinates (see section 2.3).
  • when $x_{k_{i}}-x_{k_{i}+1}=c$ for some $1\leq k_{1}<\dots<k_{\ell}\leq n$, then the level surface of $\left\{H_{k}\right\}_{k=2}^{n}$ is $(n-\ell-1)$-dimensional (the generic case corresponds to $\ell=0$), with the angle variables described in section 3.
Generic $(n-1)$-dimensional tori are well known and completely described long time ago [7]. When $\ell=n-1$, the level surface is zero-dimensional, and is a fixed point of the multitime evolution (see also [LRS]).
$x_{2}-x_{3}$$x_{1}-x_{2}$cc2222-dim1111-dim1111-dim00-dim
Figure 1. Dimensions of Liouville tori over shifted principal Weyl chamber in the $x$-plane for $G=SU(3)$.
This paper was inspired and motivated by the work [5] on hybrid integrable systems and on the upcoming paper [2] on stratified integrable systems.
Similar stratification exists in many other models, for example, in the elliptic CM system and in integrable systems on moduli spaces of flat connections. We will address this in a separate publication. Superintegrable systems and, in particular, superintegrable systems of spin CMS type are studied in [2].
@@ Acknowledgment
The work of A. L. was supported by the Beijing Natural Science Foundation (IS24006). The work of N. R. was supported by the Collaboration Grant ”Categorical Symmetries” from the Simons Foundation, by the Changjiang fund, and by the project 075-15-2024-631 funded by the Ministry of Science and Higher Education of the RF.
2. Action-angle variables in the Calogero–Moser–Sutherland system
2.1. The Calogero–Moser–Sutherland system for $SU(n)$.
Here we review well known facts about the Calogero–Moser (CM) system for $SU(n)$. This section is a well-known extension of [4] from $T^{*}\mathfrak{g}$ to $T^{*}G$. We included it for clarity of exposition. Elements of $\mathfrak{su}(n)$ and traceless $n\times n$ hermitian matrices and elements of $SU(n)$ are unimodular $n\times n$ unitary matrices.
The phase space for the Calogero–Moser system for $G=SU(n)$(and of its spin analogs) can be realized as the symplectic reduction of $T^{*}G\simeq\mathfrak{g}\times G$ with respect to the Hamiltonian action of $G$ by conjugations. From now we identify $\mathfrak{g}\simeq\mathfrak{g}^{*}$ via the Killing form. For $\mathfrak{su}(n)$ this means that both $\mathfrak{su}(n)$ and $\mathfrak{su}(n)^{*}$ are represented by traceless $n\times n$ matrices. If $\mathcal{O}$ is a coadjoint orbit in $\mathfrak{su}(n)$ of rank one, the corresponding Hamiltonian reduction is
(4) $\mathcal{S}(\mathcal{O})=\{(x,g)\in\mathfrak{su}(n)\times SU(n)\mid x-g^{-1}xg\in\mathcal{O}\}/G,$
In general, it is a stratified symplectic space. When $\mathcal{O}$ is a rank $1$ orbit, it is isomorphic to $T^{*}H_{\mathrm{reg}}/S_{n},$ where $H_{\mathrm{reg}}=\{\mathrm{diag}(\gamma_{1},\ldots,\gamma_{n});\ \gamma_{i}\neq\gamma_{j}\}$ and is a symplectic orbifold. Since the Hamiltonians are also defined on a smooth manifold $T^{*}H_{reg}$, it is convenient to work there. We will denote it $\widetilde{\mathcal{S}(\mathcal{O})}$.
A rank $1$ orbit in $\mathfrak{su}(n)$ consists of traceless Hermitian matrices of rank $1$. They can be parametrized as
$\mu_{ij}=\overline{\varphi_{i}}\varphi_{j}-c\delta_{ij},\quad c=\frac{1}{n}\sum_{i=1}^{n}|\varphi_{i}|^{2}$
modulo transformations $\varphi_{j}\mapsto e^{i\theta}\varphi_{j}$. It is clear that the second Casimir function 111 The values of higher Casimir functions are the rank $1$ orbit are determinated by $c$ and (5) $\mathrm{Tr}(\mu^{k})=c^{k}\,(n-1)\left((n-1)^{k-1}+(-1)^{k}\right).$ is
$\mathrm{tr}(\mu^{2})=c^{2}n(n-1).$
The moment map brings the $SU(n)$-orbit through $(x,g)\in T^{*}G$ to
@@ 2. Action-angle variables in the Calogero–Moser–Sutherland system
From the construction of the phase space as a Hamiltonian reduction of $T^{*}SU(n)$ with respect to rank $1$ orbit the functions
(7) $H_{k}=\frac{1}{k}\mathrm{Tr}\big{(}x^{k}\big{)},\ \ k=2,\dots,n$
on $T^{*}G=\{(x,g)\in\mathfrak{su}(n)\times SU(n)$ Poisson commute. This follows immediately from the symplectic structure on $T^{*}G=\{(x,g)\in\mathfrak{su}(n)\times SU(n)$. As a consequence the Poisson commute on $\mathcal{S}(\mathcal{O})$. Functions $\{H_{k}\}_{k=2}^{n}$ are independent on $S(\mathcal{O})$ and because $\mathrm{dim}\,S(\mathcal{O})=2n-2$ they define a Liouville integrable system.
@@ 2.2. The diagonalization of $g$ and natural coordinates
Let us briefly recall the Hamiltonian reduction construction of Calogero-Moser system by Kazhdan, Kostant, and Sternberg [4], extended to Calogero-Moser-Sutherland. We can choose a representative of the $SU(n)$–orbit through $(x,g)$ in (4) such that $g$ is diagonal:
@@
Note that if $\gamma_{i}=\gamma_{j}$, the equation (9) implies $\overline{\varphi}_{i}\varphi_{j}=0$ but this contradicts to $\lvert\varphi_{i}\rvert^{2}=c$. Thus $\gamma_{i}\neq\gamma_{j}$.
When $\gamma_{i}\neq\gamma_{j}$ the stabilizer of (8) is the Cartan subgroup of $SU(n)$ of unitary diagonal unimodular matrices. By the action of this subgroup on $x$ we can bring $\varphi_{j}$ to the form $\varphi_{j}=\sqrt{c}\,e^{i\theta}$. This gives the description of the phase space as the quotient of $T^{*}H_{reg}$ (with coordinates $(x_{ii},\gamma_{i})$) with respect to the action of $S_{n}$ by permutation of pairs $(x_{ii},\gamma_{i})$.
@@
The unimodularity implies $\sum_{i=1}^{n}p_{i}=0$. For the first two Hamiltonians we have:
@@
Here $\gamma_{j}=\exp(iq_{j})$.
2.3. The diagonalization of $x$ and the action-angle variables
Geometrically, any integrable system is a projection from a $2n$-dimensional symplectic manifold to an $n$-dimensional space, such that the generic fiber is Lagrangian.
For Calogero–Moser systems we have the projection
$\pi\colon T^{*}H_{\mathrm{reg}}\to\mathfrak{h}^{*}/W,\quad(x,\gamma)\mapsto[x]\in\mathfrak{g}^{*}/G\simeq\mathfrak{h}^{*}/W.$
Here $x$ is given by (12) as the function of $p$ and $\gamma$, and for $\mathfrak{su}(n)$. $W=S_{n}$, and $[x]$ is a collection of eigenvalues of $x$. The quotient space $\mathfrak{h}^{*}/W$ can be modeled as:
$\mathfrak{h}^{*}/W\simeq\left\{x\in\mathbb{R}^{n}\ \middle|\ x_{1}+\dots+x_{n}=0,\ x_{1}\geq\dots\geq x_{n}\right\}.$
The main observation of this paper is that the image of this mapping is a shifted cone in $\mathbb{R}^{n-1}$. Preimages of strata of this cone stratify $T^{*}H_{\mathrm{reg}}$. For CMS model we describe the strata later in this section.
2.3.1. The generic stratum
Any $x\in\mathfrak{su}(n)^{*}$ is diagonalizable by the conjugation action of $SU(n)$. let us choose a representative $(x,g)$ in our phase space where $x$ is diagonal:
(15) $x=\mathrm{diag}\left(x_{1},\ldots,x_{n}\right)\in\mathfrak{h}^{*}\subset\mathfrak{su}(n)^{*},\qquad x_{1}+x_{2}+\ldots+x_{n}=0,\qquad x_{1}\geq x_{2}\geq\dots\geq x_{n}.$
The Hamiltonians, evaluated in terms of eigenvalues of $x$ are simply symmetric power functions:
$H_{k}=\sum_{j=1}^{n}x_{j}^{k}.$
@@ 2.3.1. The generic stratum
Solutions to the momentum map equation
(16) $x-g^{-1}xg=\bar{\psi}\psi^{t}-c\cdot 1.$
can be described in the following way. Let’s $\lambda=g\bar{\psi}$, then the following holds
1) The following conditions are equivalen
@@
1) The equation for the moment map can be written as
(19) $g^{-1}xg=x+c\cdot 1-\bar{\psi}\psi^{t},$
Following lemma 3.4 from [3], let’s consider the equality of characteristic polynomials of both parts of the equation (31)
(20) $\prod_{j=1}^{n}(x_{j}-z)=\mathrm{det}(x+c\cdot 1-\bar{\psi}\psi^{t}-z)=\prod_{j=1}^{n}(x_{j}+c-z)-\sum_{j=1}^{n}|\psi_{j}|^{2}\prod_{k\neq j}(x_{k}+c-z).$
Here we take into account that for diagonal matrix $A$
(21) $\mathrm{det}\left(A+uv^{t}\right)=\mathrm{det}\left(A\right)+v^{t}A^{adj}u,$
where $A^{adj}$ is the adjugate matrix to $A$, which defined by $A^{adj}A=\mathrm{det}A\cdot 1$. This is a well-known fact which follows from Lemma 1 in the appendix A.
Equation (20) at the point $z=x_{i}+c$ is
@@
It follows from the equation (22) for $i=1$ that $|\psi_{1}|$ is strictly positive
@@
In equation (22) for $i=2$
(24) $c\ \frac{(x_{2}-x_{1}+c)}{x_{2}-x_{1}}\prod_{j>2}\frac{(x_{2}-x_{j}+c)}{(x_{2}-x_{j})}=|\psi_{2}|^{2}$
the product over $j>2$ is strictly positive. So we have
(25) $|\psi_{2}|^{2}\geq 0\qquad\Leftrightarrow\qquad x_{1}\geq x_{2}+c,$
where equality $x_{1}=x_{2}+c$ is achieved if and only if $|\psi_{2}|^{2}=0$.
Continuing this reasoning, we get the statement of the proposition. So, we have
(26) $|\psi_{k}|^{2}\geq 0\qquad\Leftrightarrow\qquad x_{k-1}\geq x_{k}+c,$
where equality $x_{k-1}=x_{k}+c$ is achieved if $|\psi_{k}|^{2}=0$.
@@
(27) $gxg^{-1}=x-c\cdot 1+\lambda\bar{\lambda}^{t},$
Equality of characteristic polynomials gives
(28) $\prod_{j=1}^{n}(x_{j}-z)=\mathrm{det}(x-c\cdot 1+\lambda\bar{\lambda}^{t}-z)=\prod_{j=1}^{n}(x_{j}-c-z)+\sum_{j=1}^{n}|\lambda_{j}|^{2}\prod_{k\neq j}(x_{k}-c-z).$
Equation (28) at the point $z=x_{i}-c$ is
@@
(30) $|\lambda_{k-1}|^{2}\geq 0\qquad\Leftrightarrow\qquad x_{k-1}\geq x_{k}+c,$
where equality $x_{k-1}=x_{k}+c$ is achieved if $|\lambda_{k-1}|^{2}=0$.
2) This part of the statement directly follows from (22) and (29).
3) After the multiplication of the momentum map (31) by $g$ the equation for the moment map becomes
(31) $gx+cg-xg=\lambda\psi^{t},$
$(i,j)$ whose element has the form
@@
which fix the form of the thirt part of the statement
which implies (18).
Remark 1.

### v2

# Low–dimensional tori in Calogero–Moser–Sutherland systems

@@ Abstract.
The main result of this paper is an explicit description of the stratification of the phase space of Calogero-Moser-Sutherland integrable systems corresponding to Lie groups $SU(n)$. Each stratum is isomorphic to $\mathbb{T}^{s}\times\mathbb{R}^{s}$ where $s=0,1,\dots,n-1$. Tori in such strata are Liouville tori. The zero-dimensional stratum is the fixed point of the multi-time dynamics of the model.
@@ 1. Introduction
Recall that Calogero–Moser–Sutherland (CMS) system [1, 6, 8] is a model of $n$ one–dimensional particles on a circle. The phase space of this model is $T^{*}H_{\rm reg}$ where $H_{\rm reg}$ is the regular part of the Cartan subgroup of $SU(n)$, i.e. diagonal unitary matrices with distinct eigenvalues. The space $T^{*}H_{\rm reg}$ has natural coordinates $(p_{i},\gamma_{i})$, $p_{i}\in\mathbb{R}$, $\lvert\gamma_{i}\rvert=1$, $\sum_{i=1}^{n}p_{i}=0$, $\prod_{i=1}^{n}\gamma_{i}=1$, $\gamma_{i}\neq\gamma_{j}$.
The standard symplectic structure on $T^{*}H_{\rm reg}$ gives the following Poisson brackets:
@@
(3) $H_{2}=\frac{1}{2}\sum_{i=1}^{n}p_{i}^{2}+\frac{1}{4}\sum_{i\neq j}^{n}\frac{c^{2}}{\sin^{2}\left(\tfrac{q_{i}-q_{j}}{2}\right)}.$
@@
where the matrix $x$ as a function of $p$ and $q$ is given by formula (18). The matrix $x$ is known as the Lax matrix for the CMS system. Functions $\{H_{k}\}_{k=2}^{n}$ are independent [^1] They are independent in the usual sense: $dH_{2}\wedge\dots\wedge dH_{n}$ is not identically zero. and Poisson commute. This follows immediately from the Hamiltonian reduction construction [4]. Functions $H_{k}$ are invariant with respect to simultaneous permutations of $p$ and $q$. Thus, the effective phase space is a manifold $S(\mathcal{O})=T^{*}H_{reg}/S_{n}$ [^2] The notation $S(\mathcal{O})$ for the phase space is natural from the point of view of the Hamiltonian reduction, see section 2.1, where $S_{n}$ acts on $\mathfrak{h}^{*}\times H_{reg}$ diagonally.
Geometrically, the CMS system is a Lagrangian fibration $\pi:S(\mathcal{O})\to\mathfrak{h}/S_{n}$. Here $\mathfrak{h}\simeq\mathbb{R}^{n-1}$ is the Cartan subalgebra of $su(n)$ and $\pi([(x,g)])=[x]$, where $[(x,g)]$ and $[x]$ are conjugacy classes of $(x,g)$ and of $x$ respectively, see section 2.1 for details. One of the main results of this paper is an explicit description of $B(\mathcal{O})=Im(\pi)$ as a stratified space which is the shifted principal Weyl chamber for $su(n)$:
(5) $B(\mathcal{O})=\sqcup_{\{j\}}B(\mathcal{O})_{\{j\}}.$
Here $\{j\}=\{j_{0},j_{1},\dots,j_{s}\}$, where $j_{0}=1<j_{1}<\dots<j_{s}$, and $0\leq s\leq n-1$. The stratum $B(\mathcal{O})_{\{j\}}$ is isomorphic to $\mathbb{R}^{s}$. On this stratum $x_{i}-x_{i+1}=c$ if $i$ is not in the set $\{j\}$ and $x_{i}-x_{i+1}>c$ if $i\in\{j\}$. The maximum stratum $B(\mathcal{O})_{\{1,2,\dots,n\}}$ has dimension n1n-1 and is the interior of the shifted principal Weyl chamber, where $x_{i}-x_{i+1}>c$ for all $i=1,\dots,n-1$. Other strata are its walls.
As a consequence the phase space $S(\mathcal{O})$ gets a stratification by the preimages of $B(\mathcal{O})_{\{j\}}$.
(6) $S(\mathcal{O})=\sqcup_{\{j\}}S(\mathcal{O})_{\{j\}},$
where $S(\mathcal{O})_{\{j\}}=\pi^{-1}(B(\mathcal{O})_{\{j\}})$. In section 3.3 we prove that
(7) $S(\mathcal{O})_{\{j\}}\simeq\mathbb{T}^{s}\times\mathbb{R}^{s}.$
The maximal stratum $S(\mathcal{O})_{\{1,2,\dots,n\}}$ is subject to the Arnold–Liouville Theorem and its structure is well known, see for example [7]. The action-angle variables on this stratum are described in section 2.3. Action-angle variables on low–dimensional strata are described in section 3.3.
The minimal stratum $S(\mathcal{O})_{\{1\}}$ is the fixed point of the multi-time Hamiltonian flow generated by Poisson commuting Hamiltonians $H_{k}$ (see [5]).
$x_{2}-x_{3}$$x_{1}-x_{2}$cc22-dim11-dim11-dim0-dim
Figure 1. The shifted principal Weyl chamber for $G=SU(3)$ and dimensions of Liouville tori over the corresponding strata.
This paper is closely related to the work [5] on hybrid integrable systems and to the upcoming paper [2] on stratified super-integrable systems.
Similar stratification exists in many other models, for example, in the elliptic CMS system and integrable systems on moduli spaces of flat connections. We will address this in a separate publication. We also do not discuss here the algebro-geometric setting.
@@ Acknowledgment
The work of A. L. was supported by the Beijing Natural Science Foundation (IS24006) and Beijing Talent Program. The work of N. R. was supported by the Collaboration Grant ”Categorical Symmetries” from the Simons Foundation, by the Changjiang fund, and by the project 075-15-2024-631 funded by the Ministry of Science and Higher Education of the Russian Federation.
2. The Calogero–Moser–Sutherland system
2.1. The Hamiltonian reduction
Here we review some facts about the CMS system for $SU(n)$. From now on, $G$ is always $SU(n)$. This section is included for completeness of the exposition. It is a well-known $T^{*}G$ version of the Kazhdan-Kostant-Sternberg construction [4] of the Calogero-Moser system by the Hamiltonian reduction of $T^{*}\mathfrak{g}$.
Here and below $\mathfrak{g}=su(n)$. We treat elements of $su(n)$ as traceless $n\times n$ Hermitian matrices, and elements of $SU(n)$ as $n\times n$ unimodular unitary matrices.
The phase space for the CMS system for $G=SU(n)$ can be realized as the Hamiltonian reduction of $T^{*}G\simeq\mathfrak{g^{*}}\times G$ with respect to the Hamiltonian action of $G$ by conjugations. From now on, we identify $\mathfrak{g}\simeq\mathfrak{g}^{*}$ via the Killing form. For $su(n)$ this means that both $su(n)$ and $su(n)^{*}$ are represented by traceless Hermitian $n\times n$ matrices.
The Hamiltonian reduction for the adjoint action of $SU(n)$ on $T^{*}SU(n)$ via the coajoint orbit $\mathcal{O}\subset su(n)^{*}$ produces the symplectic leaf of the Poisson space $T^{*}SU(n)/SU(n)$ which is
(8) $\mathcal{S}(\mathcal{O})=\mu^{-1}(\mathcal{O})/G=\{(x,g)\in su(n)\times SU(n)\mid x-g^{-1}xg\in\mathcal{O}\}/G.$
Here $\mu:T^{*}G\to\mathfrak{g}^{*}$ is the moment map for the adjoint action, $\mu(x,g)=x-g^{-1}xg$.
When $\mathcal{O}$ is a coadjoint orbit of rank one, the corresponding Hamiltonian reduction is isomorphic to $T^{*}H_{\mathrm{reg}}/S_{n}$ as a symplectic manifold. Here $H_{\mathrm{reg}}=\{\mathrm{diag}(\gamma_{1},\ldots,\gamma_{n});\ \gamma_{i}\neq\gamma_{j}\}$ [^3] Clearly $T^{*}H_{\mathrm{reg}}/S_{n}$ is a smooth symplectic manifold..
A rank $1$ orbit in $su(n)$ consists of traceless Hermitian matrices of rank $1$. They can be parametrized as
(9) $\mu_{ij}=\overline{\varphi_{i}}\varphi_{j}-c\delta_{ij},\quad c=\frac{1}{n}\sum_{i=1}^{n}|\varphi_{i}|^{2}$
modulo transformations $\varphi_{j}\mapsto e^{i\theta}\varphi_{j}$. It is clear that the second Casimir function  [^4] The values of higher Casimir functions are the rank $1$ orbit are determinated by $c$ and (10) $\mathrm{Tr}(\mu^{k})=c^{k}\,(n-1)\left((n-1)^{k-1}+(-1)^{k}\right).$ is
(11) $\mathrm{tr}(\mu^{2})=c^{2}n(n-1).$
The moment map for the adjoint action of $G$ on brings $(x,g)\in T^{*}G$ to
@@ 2.1. The Hamiltonian reduction
Functions
(13) $H_{k}(x,g)=\frac{1}{k}\mathrm{Tr}\big(x^{k}\big),\ \ k=2,\dots,n$
on $T^{*}G$ Poisson commute. This follows immediately from the symplectic structure on $T^{*}G$. Because they are $G$-invariant, they are functions on $T^{*}G/G$. As a consequence, their restriction to $S(\mathcal{O})=\mu^{-1}(\mathcal{O})/G\subset T^{*}G/G$ also Poisson commute. Functions $\{H_{k}\}_{k=2}^{n}$ are independent on $S(\mathcal{O})$ and because $\mathrm{dim}\,S(\mathcal{O})=2n-2$ they define a Liouville integrable system.
@@ 2.2. The diagonalization of $g$ and natural coordinates
Let us recall how to obtain the Hamiltonians of the CMS system.
Choose a representative of the SU(n)SU(n)–orbit through $(x,g)$ in (8) such that $g$ is diagonal:
@@
Note that if $\gamma_{i}=\gamma_{j}$, the equation (15) implies $\overline{\varphi}_{i}\varphi_{j}=0$ but this contradicts to $\lvert\varphi_{i}\rvert^{2}=c$. Thus, $\gamma_{i}\neq\gamma_{j}$.
When $\gamma_{i}\neq\gamma_{j}$ the stabilizer of (14) is the Cartan subgroup of $SU(n)$, i.e. the subgroup of unitary diagonal unimodular matrices. By the action of this subgroup on $x$ we can bring $\varphi_{j}$ to the form $\varphi_{j}=\sqrt{c}\,e^{i\theta}$. This describes the phase space as the quotient of $T^{*}H_{reg}$ (with coordinates $x_{ii}$ on $\mathfrak{g}^{*}$, and $\gamma_{i}$ on $H_{reg}\subset G$) with respect to the action of $S_{n}$ by permutation of pairs $(x_{ii},\gamma_{i})$.
@@
The tracelessness of $x$ implies $\sum_{i=1}^{n}p_{i}=0$, and the unimodularity of $\gamma$ implies $\gamma_{1}\dots\gamma_{n}=1$. For the first two Hamiltonians, we have:
@@
Here $\gamma_{j}=\exp(iq_{j})$. The Hamiltonian reduction and the Poisson structure on $T^{*}G$ imply that $G$-invariant functions on $T^{*}G\simeq\mathfrak{g}^{*}\times G$ that come from $G$-invariant functions on $\mathfrak{g}^{*}$ form a Poisson commutative subalgebra.
2.3. The CMS system as a Lagrangian fibration
Geometrically, any integrable system is a projection from a 2n2n–dimensional symplectic manifold to an nn–dimensional space, such that the generic fiber is Lagrangian. For the CMS system this projection is
(21) $\pi:\mathcal{S}(\mathcal{O})\to{su(n)^{*}}/SU(n)\simeq\mathfrak{h}^{*}/W.$
Here $\mathcal{S}(\mathcal{O})=\mu^{-1}(\mathcal{O})/G$ is the phase space obtained by the Hamiltonian reduction, and $\mathfrak{h}\subset{su(n)}$ is the Cartan subgroup (real traceless diagonal matrices), and the Weyl group $W$ is the symmetric group $S_{n}$. From the results of the previous section, the CMS Lagrangian projection is equivalent to
(22) $\pi\colon T^{*}H_{\mathrm{reg}}\to\mathfrak{h}^{*}/W,\quad(p,\gamma)\mapsto[x]\in\mathfrak{g}^{*}/G\simeq\mathfrak{h}^{*}/W.$
Here $x$, as the function of $p$ and $\gamma$, is given by (18). For $su(n)$ we can identify $[x]$ with an ordered collection $(x_{1}\geq x_{2}\geq\dots\geq x_{n})$ of eigenvalues of $x$:
(23) $\mathfrak{h}^{*}/W\simeq\left\{x\in\mathbb{R}^{n}\ \middle|\ x_{1}+\dots+x_{n}=0,\ x_{1}\geq\dots\geq x_{n}\right\}\subset\mathbb{R}^{n-1}.$
This is the principal Weyl chamber for $su(n)$. Now let us describe the image of this projection as a shifted principal Weyl chamber in $\mathbb{R}^{n-1}$ and the stratification of $T^{*}H_{\mathrm{reg}}$ by the primages of walls of the shifted principal Weyl chamber.
3. The base of the Lagrangian projection for the CMS model and the stratification of $S(\mathcal{O})$
3.1. Solutions to the moment map equation
Here we will describe the image $B(\mathcal{O})\subset\mathfrak{h}^{*}/W$ of the Lagrangian projection (22). We show that it is a convex set which can be regarded as a stratified space, which is a shifted principal Weyl chamber. We also describe the induced stratification of the phase space $S(\mathcal{O})$.
Any $x\in su(n)^{*}$ is diagonalizable by the conjugation action of $SU(n)$. Let us choose a representative $(x,g)\in\mathfrak{g}^{*}\times G$ of the point $G(x,g)\in\mu^{-1}(\mathcal{O})/G\subset T^{*}G/G$ in the phase space where $x$ is diagonal:
(24) $x=\mathrm{diag}\left(x_{1},\ldots,x_{n}\right)\in\mathfrak{h}^{*}\subset su(n)^{*},\qquad x_{1}+x_{2}+\ldots+x_{n}=0,\qquad x_{1}\geq x_{2}\geq\dots\geq x_{n}.$
@@ 3.1. Solutions to the moment map equation
Solutions to the momentum map equation [^5] From now on, when it is not ambiguous, we will write $a$ instead of $a\cdot\text{Id}$ for the matrix which is a multiple of the identity matrix.
(25) $x-g^{-1}xg=\bar{\psi}\psi^{t}-c\ \$
can be described in the following way. Define $\lambda=g\bar{\psi}$, then the following holds:
1) The following conditions are equivalent
@@
Let us start with part 1). The equation for the moment map can be written as
(28) $g^{-1}xg=x+c-\overline{\psi}\psi^{t}.$
Let us consider the equality of characteristic polynomials of both parts of this equation [^6] Compare with Lemma 3.4 from [3].,
(29) $\prod_{j=1}^{n}(x_{j}-z)=\mathrm{det}(x+c-\bar{\psi}\psi^{t}-z)=\prod_{j=1}^{n}(x_{j}+c-z)-\sum_{j=1}^{n}|\psi_{j}|^{2}\prod_{k\neq j}(x_{k}+c-z).$
Here we take into account that for a diagonal matrix AA
(30) $\mathrm{det}\left(A+uv^{t}\right)=\mathrm{det}\left(A\right)+v^{t}A^{\text{adj}}u,$
where $A^{\text{adj}}$ is the adjugate matrix to $A$, which defined by $A^{\text{adj}}A=\mathrm{det}A\cdot{\rm Id}$. This is a well-known fact which follows from lemma 1 in appendix A.
The identity (29) for $z=x_{i}+c$ is
@@
which proves (26).
The equation (31) for $i=1$ implies that $|\psi_{1}|$ is strictly positive
@@
In the equation (31) for $i=2$ the product over $j>2$ is strictly positive:
[…]
This implies that
(34) $|\psi_{2}|^{2}\geq 0$
if and only if $x_{1}\geq x_{2}+c$ and the equality $x_{1}=x_{2}+c$ is achieved if and only if $|\psi_{2}|^{2}=0$.
Continuing this reasoning, we get:
(35) $|\psi_{k}|^{2}\geq 0\ \$
if and only if $x_{k-1}\geq x_{k}+c$ and the equality $x_{k-1}=x_{k}+c$ is achieved only if $|\psi_{k}|^{2}=0$.
@@
(36) $gxg^{-1}=x-c+\lambda\bar{\lambda}^{t}.$
The equality of characteristic polynomials gives
(37) $\prod_{j=1}^{n}(x_{j}-z)=\mathrm{det}(x-c+\lambda\bar{\lambda}^{t}-z)=\prod_{j=1}^{n}(x_{j}-c-z)+\sum_{j=1}^{n}|\lambda_{j}|^{2}\prod_{k\neq j}(x_{k}-c-z).$
Evaluating this idenity (37) at the point $z=x_{i}-c$ gives
@@
(39) $|\lambda_{k-1}|^{2}\geq 0$
if and only if $x_{k-1}\geq x_{k}+c$, where equality $x_{k-1}=x_{k}+c$ is achieved only if $|\lambda_{k-1}|^{2}=0$. This proves part 1).
Part 2) directly follows from (31) and (38). We already proved 3).
Let us prove 4). After the multiplication of the momentum map (40) by $g$, the equation for the moment map becomes
(40) $gx+cg-xg=\lambda\psi^{t}\ \ .$
Matrix elements of this equation are a system of equations for $g_{ij}$:
@@
which implies (27). ∎
Corollary 1.
The image of the projection (22) is the subset $B(\mathcal{O})\subset\mathbb{R}^{n-1}/S_{n}$ which can be described as
(42) $B(\mathcal{O})=\{(x_{1},\dots,x_{n})\mid x_{i}-x_{i+1}\geq c,\ \sum_{i=1}^{n}x_{i}=0\}.$
This space is naturally stratified. Let $\{j\}=\{j_{0},j_{1},\dots,j_{s}\}$ be a collection where $j_{0}=1,1<j_{a}<j_{a+1}\leq n-1$. Then
(43) $B(\mathcal{O})=\sqcup_{\{j\}}B(\mathcal{O})_{\{j\}},$
where
[…]
It is easy to check that
(45) $\overline{B(\mathcal{O})_{\{j\}}}=\sqcup_{\{k\}\leq\{j\}}B(\mathcal{O})_{\{k\}}.$
Here $\{k\}\leq\{j\}$ means that $\{k\}$ is a subset of $\{j\}$. This stratification for $G=SU(3)$ is illustrated on Figure 1.
This stratification of $B(\mathcal{O})$ induces a stratification of $S(\mathcal{O})$ by the primages of $B$-strata, which we will describe below.
3.2. The maximal stratum